OpenStudy (anonymous):

how do you solve the proportion 15t/5=2t+3/6

OpenStudy (lucaz):

subtract 2t from both sides

OpenStudy (anonymous):

\[\frac{ 15t }{ 5}= \frac{ 2t+3 }{ 6 } this is how \it looks\]

OpenStudy (anonymous):

this is how it looks *

OpenStudy (lucaz):

oh, okay..

OpenStudy (lucaz):

multiply both sides by 6 and 5 to get rid off the denominators

OpenStudy (lucaz):

or just sinplify 15t/5 = 3t*5/5 = 3t

OpenStudy (lucaz):